September 3, 2015 Drift Chamber meeting
- Instructions for Bluejeans meeting connection
- Meeting ID: 290664653
- To join via a Web Browser, go to the page  https://bluejeans.com/290664653.
- CDC update (Mike, Simon, Beni)
- FDC update (Luke,Lubomir)
Participants: Curtis, Naomi (CMU), Fernando, Eugene, Luke, Beni, Dave, Marc (Mcmullen), Chris, Nick, Simon, Sergey and Lubomir (JLab).
- Simon went trough half of the CDC layers: the deformations are not always in vertical direction, i.e. this may not be sagging. Need to make a map of all the measurements to make some conclusion.
- Beni: CDC is working now with the new MFC (separated from the FDC) blending the same 50/50 mixture. Need to verify this with the CO2 sensor.
- Measurements over weekend in mode 7.
- Lubomir showed cathode resolution results from the spring run with two different ST+FCAL triggers: run 3137 "standard" trigger and run 3200 excluding 4 additional layers of blocks in the center. The resolutions with the inner blocks excluded looks better and now very close to the one (~200 microns) in the fall run. This was achieved by using tighter cuts on the tracks used for the resolution studies; obviously the fact that we had half the energy in the spring explains this. Still, the offsets used for the fall and the spring runs are different and we don't understand why.
- Studies with 55Fe on the thick cell in the spare package: due to diffusion not all the charge is visible depending how far the photon interaction was from the wire. Expecting better results with 90/10 gas mixture where the drift times are much shorter.
- The reason for the LV module failure not known yet.
- Dave is planing to shut down the interlock systems on FDC and CDC whenever that is convenient.